# NoisyRollout: Reinforcing Visual Reasoning with Data Augmentation

### v1

@@ Abstract
Recent advances in reinforcement learning (RL) have strengthened the reasoning capabilities of vision-language models (VLMs). However, enhancing policy exploration to more effectively scale test-time compute remains underexplored in VLMs. In addition, VLMs continue to struggle with imperfect visual perception, which in turn affects the subsequent reasoning process. To this end, we propose NoisyRollout, a simple yet effective RL approach that mixes trajectories from both clean and moderately distorted images to introduce targeted diversity in visual perception and the resulting reasoning patterns. Without additional training cost, NoisyRollout enhances the exploration capabilities of VLMs by incorporating a vision-oriented inductive bias. Furthermore, NoisyRollout employs a noise annealing schedule that gradually reduces distortion strength over training, ensuring benefit from noisy signals early while maintaining training stability and scalability in later stages. With just $\mathbf{2.1}$K training samples, NoisyRollout achieves state-of-the-art performance among open-source RL-tuned models on $\mathbf{5}$ out-of-domain benchmarks spanning both reasoning and perception tasks, while preserving comparable or even better in-domain performance.
Figure 1: Accuracy improvement over the baseline model (Qwen2.5-VL-7B-Instruct) on out-of-domain benchmarks covering both reasoning and perception tasks. Both Qwen2.5-VL-GRPO-7B and NoisyRollout-7B are fine-tuned by ourselves (denoted with $\ddagger$) using GRPO with $2.1$K training samples from Geometry3K. The exact accuracy of NoisyRollout-7B is annotated above each corresponding bar in parentheses.
@@ 1 Introduction
Test-time compute scaling (often referred to as reasoning) through reinforcement learning (RL) has emerged as a promising axis for scaling model intelligence (Jaech et al., 2024; Cui et al., 2025). While this idea has been primarily explored in the context of large language models (LLMs), the vision-language model (VLM) community is also actively investigating this direction (Liu et al., 2025b; Meng et al., 2025). As natural extensions of LLMs for visual understanding, VLMs could intuitively benefit from similar scaling behaviors via RL (Huang et al., 2025; Wang et al., 2025a; Lu et al., 2025; Yu et al., 2025a).
@@
However, scaling test-time compute through RL requires more than sheerly generating longer outputs (Liu et al., 2025a). A key challenge is effective policy exploration, which enables models to discover behaviors that generalize well beyond training data (Yu et al., 2025b), which remains largely underexplored in VLM studies. Traditional practices, such as increasing rollout temperature to promote decoding diversity (Zeng et al., 2024), often introduce superficial variability without meaningfully guiding the policy toward more informative behaviors. Meanwhile, VLMs frequently struggle with imperfect visual perception due to the inherent complexity of visual information (Liu et al., 2024a; Wei et al., 2024). These perceptual limitations result in a significant bottleneck, especially for reasoning tasks that seriously depend on accurate interpretations of visual inputs (Zhang et al., 2024a; Jiang et al., 2025).
Tackling the challenges of policy exploration and perceptual limitations in VLMs, we propose NoisyRollout, a simple yet effective RL fine-tuning approach for VLMs that introduces meaningful rollout diversity by incorporating noisy trajectories generated from image-text pairs with distorted visual inputs, enabling policy models to perform more targeted exploration during training. The core process happpens at the rollout collection phase: the old policy $(\pi_{\theta_{\mathrm{old}}})$ generates reasoning trajectories using both original clean images and moderately distorted versions of the same images in response to identical text queries. While the current policy $(\pi_{\theta})$ is updated using only clean image inputs, trajectories from both clean and distorted images contribute to calculating the reward baseline and normalized advantage in Group Relative Policy Optimization (GRPO) (Shao et al., 2024). [^1] We adopt GRPO due to its natural compatibility with our method; demonstrating generality across other RL objectives is left for future work.
This hybrid rollout strategy strengthens learning through two key mechanisms. First, successful reasoning trajectories from distorted images reveal alternative, potentially more robust reasoning strategies, providing diverse explorations that improves generalization to harder or out-of-domain visual conditions. Second, when the same query leads to diverging outcomes between clean and distorted inputs, the resulting reward differences highlight perceptual discrepancies that affect reasoning. These discrepancies serve as implicit contrastive signals that refine the model’s perception during the reasoning process via squeezing the negative perceptual exploration space. To further improve scalability and stabilize training, we employ a noise annealing schedule that gradually reduces the strength of input distortions over training, avoiding significant distributional mismatch caused by noisy trajectories (which easily occurs in later training stages) while preserving the benefits of noisy signals in early training stages.
Extensive experiments validate the effectiveness of NoisyRollout. With just $2.1$K RL training samples from Geometry3K, NoisyRollout achieves state-of-the-art performance across $5$ challenging out-of-domain visual reasoning and perception benchmarks (Lu et al., 2023; Guan et al., 2024; Zhang et al., 2024a; Wang et al., 2024; Qiao et al., 2024) compared to other open-source RL-tuned models (Meng et al., 2025; Gui & Ren, 2025; Wang et al., 2025b) and even those leveraging large-scale SFT before RL training (Yang et al., 2025; Zhang et al., 2025a; Deng et al., 2025), with notable results including MathVerse ($53.2$%), MathVision ($28.5$%) and HallusionBench ($72.1\%$). When compared to its direct baseline (vanilla GRPO), NoisyRollout consistently outperforms it on both in-domain and out-of-domain benchmarks. The improvements generalize to other training corpora as well, showing similar gains when trained on K12 (Meng et al., 2025). Overall, NoisyRollout offers a simple and effective strategy to enhance visual reasoning with noisy reinforcement leanring, improving model generalization capabilities without extra training overhead or modifications to the underlying RL objective.
@@ 2 NoisyRollout: A Free-Lunch with Noisy Reinforcement Learning
Figure 2: Illustration of our NoisyRollout workflow. Solid lines denote the standard GRPO process, while dashed lines depict the generation and use of noisy rollouts from distorted images. KL divergence loss is omitted as discussed in Section 2. The terms $\{\mathbf{o}_{i}\}_{i=1}^{n_{1}+n_{2}}$, $\{r_{i}\}_{i=1}^{n_{1}+n_{2}}$, and $\{A_{i}\}_{i=1}^{n_{1}+n_{2}}$ represent mixed trajectories, rewards, and normalized advantages, respectively, with $n_{1}$ from clean inputs and $n_{2}$ from noisy inputs. $(I,\mathbf{q})$ and $(\tilde{I},\mathbf{q})$ denote clean and noisy inputs where $\tilde{I}=T_{\alpha_{t}}(I)$ applies a distortion with strength $\alpha_{t}$. The distortion level $\alpha_{t}$ is determined by a noise annealing schedule $\eta(\cdot)$, which gradually reduces distortion over the training process. Only clean inputs are used for policy optimization in RL.
We introduce NoisyRollout, an RL fine-tuning method designed to enhance visual reasoning in VLMs, particularly improving the rollout diversity for effective policy exploration. NoisyRollout achieves this by incorporating a hybrid rollout strategy that leverages reasoning trajectories form both clean and distorted images, and a noise annealing schedule that progressively reduces distortion strength. These adjustments require no additional training cost and can be seamlessly applied to the standard implementation of GRPO. A simplified overview is provided in Figure 2 and Algorithm 1.
GRPO. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) was originally developed to improve mathematical reasoning in LLMs but can be effectively adapted to enhance visual reasoning in VLMs. For a given input pair $(I,\mathbf{q})$ of image and text query from the training set $p_{\mathcal{D}}$, to avoid reward hacking, a rule-based reward function $R(I,\mathbf{q},\mathbf{o})$ is typically adopted such that $R(I,\mathbf{q},\mathbf{o})=1$ if the generated response $\mathbf{o}$ correctly addresses the query (as evaluated by a verifiable parser) and $R(I,\mathbf{q},\mathbf{o})=0$ otherwise. The old policy $\pi_{\theta_{\mathrm{old}}}$ generates $n$ response rollouts for this input, and the baseline reward is computed as $\bar{R}=\frac{1}{n}\sum_{i=1}^{n}R_{i}$. The normalized advantage for the $i$-th rollout is given by $\hat{A}_{i}=\frac{R_{i}-\bar{R}}{\sigma(R)}$, where $\sigma(R)$ denotes the standard deviation of rewards within the group. Derived from PPO (Schulman et al., 2017), the GRPO objective function is:
$L_{\mathrm{GRPO}}(\theta)={}$ $\mathbb{E}_{(I,\mathbf{q})\sim p_{\mathcal{D}},\mathbf{o}\sim\pi_{\theta_{\text{old}}}(\cdot|I,\mathbf{q})}$
@@
where $\pi_{\theta}$ is the current policy, $\epsilon>0$ sets the clipping range. We omit the KL divergence constraint $D_{\mathrm{KL}}[\pi_{\theta}|\pi_{\theta_{\mathrm{ref}}}]$ following recent practices in Meng et al. (2025) and Liu et al. (2025a).
Hybrid rollout strategy. Building upon GRPO, NoisyRollout introduces a hybrid rollout strategy to specifically enhance the rollout diversity from the visual perspective. The core idea is to leverage rollouts generated from both clean and distorted images, thereby enriching the signals used for policy updates. For each input pair $(I,\mathbf{q})$, we generate a corresponding noisy image $\tilde{I}$ using a noise transformation function $T$ parameterized by a noise strength $\alpha$, denoted as $\tilde{I}=T_{\alpha}(I)$. As illustrated in Figure 2, the old policy $\pi_{\theta_{\mathrm{old}}}$ then produces $n_{1}$ rollouts conditioned on the clean image $I$ and $n_{2}$ rollouts conditioned on the noisy image $\tilde{I}$. These responses are grouped together and collectively used to compute a common baseline reward, $\bar{R}=\frac{1}{n_{1}+n_{2}}\sum_{i=1}^{n_{1}+n_{2}}R_{i}$, as well as corresponding normalized advantages for each rollout. Crucially, although the noisy rollouts contribute to the baseline and advantage calculation, the policy update step remains conditioned solely on the clean image $I$ (and query $\mathbf{q}$). The NoisyRollout objective function is defined as:
$\!\!L(\theta)={}$ $\mathbb{E}_{(I,\mathbf{q})\sim p_{\mathcal{D}},\{\mathbf{o}_{j}\}_{j=1}^{n_{1}}\sim\pi_{\theta_{\mathrm{old}}}(\cdot|I,\mathbf{q}),\{\mathbf{o}_{k}\}_{k=n_{1}+1}^{n_{1}+n_{2}}\sim\pi_{\theta_{\mathrm{old}}}(\cdot|\tilde{I},\mathbf{q})}$
@@
Algorithm 1 NoisyRollout: Noisy Reinforcement Fine-Tuning
1:Input: Current policy $\pi_{\theta}$, old policy $\pi_{\theta_{old}}$, dataset $\mathcal{D}$, training steps $T_{\text{max}}$, clean rollout number $n_{1}$, noisy rollout number $n_{2}$, clip parameter $\epsilon$, initial noise strength $\alpha_{0}$, noise scheduler $\eta(\cdot)$, noise transformation function $T(\cdot)$
2:for $t=1$ to $T_{\text{max}}$ do
3:     Sample batch $(I,\mathbf{q})\sim\mathcal{D}$
5:     Generate distorted images $\tilde{I}=T_{\alpha_{t}}(I)$
9:     Compute baseline $\bar{R}=\frac{1}{n_{1}+n_{2}}\sum_{i=1}^{n_{1}+n_{2}}R_{i}$
10:     Compute advantages $\hat{A}_{i}=\frac{R_{i}-\bar{R}}{\sigma(R)}$
11:     Update policy using:
12:     $L(\theta)=\mathbb{E}\Big{[}\frac{1}{n_{1}+n_{2}}\sum_{i=1}^{n_{1}+n_{2}}\min\Big{(}\frac{\pi_{\theta}(\mathbf{o}_{i}\mid I,\mathbf{q})}{\pi_{\theta_{\text{old}}}(\mathbf{o}_{i}\mid I,\mathbf{q})}\hat{A}_{i},\mathrm{clip}\big{(}\frac{\pi_{\theta}(\mathbf{o}_{i}\mid I,\mathbf{q})}{\pi_{\theta_{\text{old}}}(\mathbf{o}_{i}\mid I,\mathbf{q})},1-\epsilon,1+\epsilon\big{)}\hat{A}_{i}\Big{)}\Big{]}$
13:     $\theta\leftarrow\theta-\nabla_{\theta}L(\theta)$triangleright Update conditioned on clean images only
14:     $\theta_{\text{old}}\leftarrow\theta$triangleright Update old policy parameters
15:end for
Noise annealing schedule. Applying fixed-strength distortions throughout training often leads to training instability, primarily due to a distributional mismatch between noisy rollouts and the evolving policy. To mitigate this, we introduce a noise annealing schedule $\eta(\cdot)$ that gradually reduces the distortion strength over time. Specifically, at training step $t$, the noise level is defined as $\alpha_{t}=\eta(\alpha_{0},t,T_{\text{max}})$, where $\alpha_{0}$ is the initial noise strength and $T_{\text{max}}$ denotes the total number of training steps. The distorted input is then generated as $\tilde{I}=T_{\alpha_{t}}(I)$. This schedule provides diverse and informative supervision signals in the early stages of training, when the policy is constrained by its perceptual capacity. As training progresses, the noise level is gradually reduced, narrowing the gap between noisy rollouts and the outputs the policy would produce on clean images. This helps prevent abrupt distribution shifts after policy updates and allows the noisy rollouts to become increasingly “on-policy” over time, contributing to more stable training dynamics and better exploitation in the later stages of training.
Summary. NoisyRollout aims to improve the visual reasoning abilities of VLMs by enhancing rollout diversity to enable more effective exploration during RL training. Built on top of GRPO, it introduces a hybrid rollout strategy and a noise annealing schedule. These additions require no extra training cost and preserve the original RL objective. This design offers several benefits:

### v3

@@ Abstract
Recent advances in reinforcement learning (RL) have strengthened the reasoning capabilities of vision-language models (VLMs). However, enhancing policy exploration to better scale test-time compute remains largely underexplored. In addition, VLMs continue to struggle with imperfect visual perception, which in turn affects the subsequent reasoning process. To this end, we propose NoisyRollout, a simple yet effective data augmentation method that mixes trajectories from both clean and moderately distorted images during RL training. By injecting targeted diversity in visual perception and the resulting reasoning patterns, NoisyRollout promotes better policy exploration through vision-oriented inductive biases, ultimately leading to more robust reasoning behaviors. We further adopt a noise annealing schedule that gradually reduces distortion strength over training, leveraging noisy signals early on while ensuring training stability in later stages. Crucially, our method is easy-to-adopt—requiring no additional training cost and no modifications to the RL objective. Extensive experiments on $\mathbf{2}$ distinct training datasets demonstrate that NoisyRollout achieves state-of-the-art performance among open-source RL-tuned models across $\mathbf{5}$ out-of-domain reasoning and perception benchmarks. Furthermore, we validate the effectiveness of NoisyRollout across model sizes ($7$B and $32$B) and data scales (from $1$K to $6$K), highlighting its generalizability and scalability.
@@ 1 Introduction
Scaling test-time compute—often referred to as reasoning—through reinforcement learning (RL) has emerged as a promising axis for advancing model intelligence [23, 11]. While this idea has been primarily explored in the context of large language models (LLMs) [20, 79], the vision-language model (VLM) community is also actively investigating this direction [41, 51, 48]. Recent endeavours suggests that VLMs can also benefit from RL-driven scaling of test-time compute [21, 39, 62, 44, 75].
@@
However, scaling test-time compute via RL requires more than sheerly generating longer outputs [40], and VLMs face unique challenges in this process. A key challenge is effective policy exploration, enabling policies to discover behaviors that generalize well beyond training data [76, 72]—an area largely underexplored in VLM research. Traditional practices, such as increasing rollout temperature to promote decoding diversity [78], often introduce superficial variability without meaningfully directing policies toward more robust or informative behaviors. Moreover, VLMs inherently struggle with imperfect visual perception [35, 68], which negatively impacts subsequent reasoning processes [82, 86, 25]. Despite this, recent efforts [41, 48, 13] tend to adapt RL methods directly from the LLM domain. Such approaches often fail to address these perceptual challenges, thereby hindering the efficient development of visual reasoning capabilities through RL.
Figure 1: Accuracy improvement over Qwen2.5-VL-7B-Instruct on $5$ out-of-domain benchmarks, covering both visual reasoning tasks (from MathVerse to WeMath) and a visual perception task (HallusionBench). Both Qwen2.5-VL-GRPO-7B and NoisyRollout-7B are fine-tuned by ourselves (denoted with $\ddagger$) using vanilla GRPO with only $\mathbf{2.1}$K training samples from Geometry3K. The exact accuracy of NoisyRollout-7B is annotated above each corresponding bar in parentheses.
Tackling the challenges of policy exploration and perceptual limitations in VLMs during RL training, we propose NoisyRollout, a simple yet powerful data augmentation technique for VLMs that introduces meaningful rollout diversity. Specifically, for each training sample consisting of an input image $I$ and a corresponding text query $\mathbf{q}$, the old policy ($\pi_{\theta_{\mathrm{old}}}$) produces two sets of rollouts based on the original clean image and a moderately distorted version of the same image, respectively. While the current policy ($\pi_{\theta}$) is updated solely by conditioning on the clean image and text query pair $(I,\mathbf{q})$, the two sets of rollouts form a group, collectively contributing to computing the reward baseline and normalized advantage in Group Relative Policy Optimization (GRPO) [56]. This hybrid rollout strategy enables the policy to achieve more targeted and efficient exploration, ultimately leading to more robust visual reasoning via RL through two key mechanisms:
  • Successful reasoning trajectories from noisy inputs with distorted images reveal alternative, potentially more robust reasoning strategies, improving reasoning generalization to harder or out-of-domain visual conditions.
  • When the same query yields different outcomes for clean and distorted inputs, the resulting reward differences expose perceptual discrepancies that affect reasoning. These discrepancies act as implicit contrastive signals, helping refine the model’s visual perception during reasoning by constraining the negative perceptual exploration space.
While incorporating noisy rollouts can facilitate more effective and efficient exploration, it may also introduce instability in policy gradient estimation. To further enhance scalability and training stability, we employ a noise annealing schedule that gradually reduces the strength of image distortions over training. Such a strategy mitigates distributional mismatch between the evolving policy and the noisy trajectories generated from it when conditioned on clean inputs—an issue that often arises in later training stages—while retaining the benefits of noisy signals during the early phases of training.
We conduct extensive experiments to validate the effectiveness of NoisyRollout. Trained with only $2.1$K samples from the Geometry3K [42] dataset using Qwen2.5-7B-VL-Instruct [4], Figure 1 shows that NoisyRollout achieves superior performance across $5$ out-of-domain visual reasoning and perception benchmarks [43, 18, 82, 64, 52] (MathVerse $53.2$%, MathVision $28.5$%, and HallusionBench $72.1\%$). It outperforms both open-source RL-tuned models [48, 67] and those utilizing large-scale supervised fine-tuning (SFT) before RL [73, 81, 13]. Furthermore, it consistently surpasses its direct baseline (vanilla GRPO) on both in-domain and out-of-domain tasks, all within a fixed total rollout budget. Crucially, these out-of-domain improvements generalize across different model sizes (e.g., $7$B to $32$B) as well as training corpora and data scales (e.g., MMK12 [48] with $1$K to $6$K samples). These empirical results, combined with its simplicity and lightweight characteristics, establish NoisyRollout as a potentially scalable approach. We summarize the contributions of this paper as follows:
  • We identify and investigate critical, yet underexplored, challenges in policy exploration and perceptual robustness that arise during the RL training of VLMs.
  • We introduce NoisyRollout, advancing a data augmentation perspective to improve visual reasoning generalization, by mixing trajectories from both clean and moderately distorted images.
  • Extensive experiments demonstrate NoisyRollout’s consistent outperformance of vanilla GRPO across model sizes and data scales, achieving state-of-the-art results among open-source RL-tuned VLMs—without additional training overhead or complex system modifications.
Figure 2: Illustration of the NoisyRollout workflow. Solid lines depict the generation and use of clean rollouts from the clean (original) input $(I,\mathbf{q})$, while dashed lines depict the generation and use of noisy rollouts from the corresponding noisy input $(\tilde{I},\mathbf{q})$. The distorted image $\tilde{I}$ is obtained by applying a distortion function $\tilde{I}=T_{\alpha_{t}}(I)$ with distortion strength $\alpha_{t}$. The distortion level $\alpha_{t}$ is controlled by a noise annealing schedule, which gradually decreases distortion during training. The terms $\{\mathbf{o}_{i}\}_{i=1}^{n_{1}+n_{2}}$, $\{r_{i}\}_{i=1}^{n_{1}+n_{2}}$, and $\{A_{i}\}_{i=1}^{n_{1}+n_{2}}$ represent mixed trajectories, rewards, and advantages, respectively. Notably, policy optimization conditions only on clean inputs, while the corresponding noisy inputs are used solely to collect noisy rollouts.
@@ 2 NoisyRollout: A Free-Lunch with Noisy Reinforcement Learning
We introduce NoisyRollout, a data augmentation method that enhances visual reasoning in VLMs during RL training, particularly by improving the rollout diversity for better policy exploration. NoisyRollout achieves this by incorporating a hybrid rollout strategy that leverages reasoning trajectories from both clean and distorted images, and a noise annealing schedule that progressively reduces distortion strength. These designs require no additional training cost and integrate seamlessly with standard GRPO implementations. A simplified overview is provided in Figure 2 and Algorithm 1.
GRPO. Group Relative Policy Optimization (GRPO) [56] was originally developed to improve mathematical reasoning in LLMs but can also be effectively adapted to enhance visual reasoning in VLMs. For a given input pair $(I,\mathbf{q})$ consisting of an image and text query from the training set $p_{\mathcal{D}}$, a rule-based outcome reward function $r(I,\mathbf{q},\mathbf{o})$ is adopted to avoid reward hacking. This function assigns $r(I,\mathbf{q},\mathbf{o})=1$ if the generated response $\mathbf{o}$ correctly addresses the query (as verified by a parser) with the required format, and $r(I,\mathbf{q},\mathbf{o})=0$ otherwise. For each input, the old policy $\pi_{\theta_{\mathrm{old}}}$ generates $n$ response rollouts. The baseline reward is then calculated as $\mathrm{mean}(\mathbf{r})$, where $\mathbf{r}=\{r_{i}\}_{i=1}^{n}=\{r(I,\mathbf{q},\mathbf{o}_{i})\}_{i=1}^{n}$ represents the rewards for all rollouts. The normalized advantage for the $i$-th rollout is defined as $\hat{A}_{i}=\frac{r_{i}-\mathrm{mean}(\mathbf{r})}{\mathrm{std}(\mathbf{r})}$. Derived from PPO [55], the GRPO objective function is:
$\mathcal{J}_{\mathrm{GRPO}}(\theta)={}$ $\mathbb{E}_{(I,\mathbf{q})\sim p_{\mathcal{D}},\mathbf{o}\sim\pi_{\theta_{\mathrm{old}}}(\cdot|I,\mathbf{q})}$
@@
where $\pi_{\theta}$ is the current policy, $\epsilon>0$ sets the clipping range. We omit the KL divergence constraint $\mathbb{D}_{\mathrm{KL}}[\pi_{\theta}|\pi_{\theta_{\mathrm{ref}}}]$ following recent practices in Meng et al. [48] and Liu et al. [40].
Hybrid rollout strategy. Building upon GRPO, NoisyRollout introduces a hybrid rollout strategy to enhance the rollout diversity. For each input pair $(I,\mathbf{q})$, we generate an augmented version of the image $\tilde{I}$ through a noise transformation function $T_{\alpha}$ parameterized by a distortion strength $\alpha$, i.e., $\tilde{I}=T_{\alpha}(I)$. As illustrated in Figure 2, the old policy $\pi_{\theta_{\mathrm{old}}}$ produces two sets of rollouts: $n_{1}$ responses conditioned on the clean input $(I,\mathbf{q})$, and $n_{2}$ responses conditioned on the corresponding noisy input $(\tilde{I},\mathbf{q})$. All rollouts from both clean and distorted images are then combined into a single group for reward calculation, yielding $\mathbf{r}=\{r_{i}\}_{i=1}^{n_{1}+n_{2}}=\{r(I,\mathbf{q},\mathbf{o}_{j})\}_{j=1}^{n_{1}}\cup\{r(I,\mathbf{q},\mathbf{o}_{k})\}_{k=n_{1}+1}^{n_{1}+n_{2}}$. Crucially, the policy update step remains conditioned solely on the clean image $I$ and query $\mathbf{q}$ for better policy exploration. We defer the discussion of optimizing noisy and clean trajectories on their corresponding inputs to Appendix B. The NoisyRollout objective function is defined as:
$\!\!\!\mathcal{J}(\theta)={}$ $\mathbb{E}_{(I,\mathbf{q})\sim p_{\mathcal{D}},\{\mathbf{o}_{j}\}_{j=1}^{n_{1}}\sim\pi_{\theta_{\mathrm{old}}}(\cdot|I,\mathbf{q}),\{\mathbf{o}_{k}\}_{k=n_{1}+1}^{n_{1}+n_{2}}\sim\pi_{\theta_{\mathrm{old}}}(\cdot|\tilde{I},\mathbf{q})}$
@@
Noise annealing schedule. Applying fixed-strength distortions throughout training often leads to training instability, primarily due to a distributional mismatch between noisy rollouts and the evolving policy. To mitigate this, we introduce a noise annealing schedule $\eta(\cdot)$ that gradually reduces the distortion strength over time. Specifically, at training step $t$, the noise level is defined as $\alpha_{t}=\eta(\alpha_{0},t,t_{\text{max}})$, where $\alpha_{0}$ is the initial noise strength and $t_{\text{max}}$ denotes the total number of training steps. As shown in Figure 2, the distorted image is then generated as $\tilde{I}=T_{\alpha_{t}}(I)$.
Consequently, this schedule keeps diverse and informative supervision signals early in training, when the policy is constrained by its perceptual capacity. As training progresses, the noise level $\alpha_{t}$ is gradually reduced, narrowing the gap between noisy rollouts ($\{\mathbf{o}_{k}\}_{k=n_{1}+1}^{n_{1}+n_{2}}$) and the trajectories that $\pi_{\theta_{\mathrm{old}}}(\cdot|I,\mathbf{q})$ would typically produce. This decay helps mitigate abrupt distribution shifts after policy updates, which can arise from unstable or high-variance policy gradients. Over time, rollouts generated from $(\tilde{I},\mathbf{q})$ become progressively more “on-policy” w.r.t the clean-input-conditioned policy $\pi_{\theta_{\mathrm{old}}}(\cdot|I,\mathbf{q})$, fostering a smoother transition from exploration to exploitation in later training stages.
Summary. NoisyRollout aims to improve the visual reasoning abilities of VLMs by enhancing rollout diversity to enable more effective policy exploration during RL training. Built on top of GRPO, it introduces a hybrid rollout strategy and a noise annealing schedule. These additions require no extra training cost and preserve the original RL objective. This design offers several benefits: